# LoRA vs Full Fine-tuning: An Illusion of Equivalence

### v1

@@ Abstract
Fine-tuning is a crucial paradigm for adapting pre-trained large language models to downstream tasks. Recently, methods like Low-Rank Adaptation (LoRA) have been shown to match the performance of fully fine-tuned models on various tasks with an extreme reduction in the number of trainable parameters. Even in settings where both methods learn similarly accurate models, are their learned solutions really equivalent? We study how different fine-tuning methods change pre-trained models by analyzing the model’s weight matrices through the lens of their spectral properties. We find that full fine-tuning and LoRA yield weight matrices whose singular value decompositions exhibit very different structure; moreover, the fine-tuned models themselves show distinct generalization behaviors when tested outside the adaptation task’s distribution. More specifically, we first show that the weight matrices trained with LoRA have new, high-ranking singular vectors, which we call intruder dimensions. Intruder dimensions do not appear during full fine-tuning. Second, we show that LoRA models with intruder dimensions, despite achieving similar performance to full fine-tuning on the target task, become worse models of the pre-training distribution and adapt less robustly to multiple tasks sequentially. Higher-rank, rank-stabilized LoRA models closely mirror full fine-tuning, even when performing on par with lower-rank LoRA models on the same tasks. These results suggest that models updated with LoRA and full fine-tuning access different parts of parameter space, even when they perform equally on the fine-tuned distribution. We conclude by examining why intruder dimensions appear in LoRA fine-tuned models, why they are undesirable, and how their effects can be minimized.
@@ 1 Introduction
Figure 1: Spectral dissimilarities between full fine-tuning and LoRA. Similarity matrix of pre- and post-fine-tuning singular vectors of the weight matrices to characterize spectral differences introduced upon fine-tuning, in a representative example for LLaMA-2 fine-tuned on Magicoder. Full fine-tuning retains most of the pre-training structure; the diagonal shift in LoRA corresponds to the introduction of intruder dimensions. Color shows cosine similarity.
@@
Adapting large, pre-trained models to downstream tasks via fine-tuning is a computation- and data-efficient way to create domain-specific models for a variety of tasks. The simplest approach is to fine-tune all parameters of the pre-trained model on downstream task data (Devlin et al., 2019; Ouyang et al., 2022). However, as pre-trained models grow larger, full fine-tuning becomes increasingly challenging and expensive. Recently, parameter-efficient fine-tuning (PEFT) methods, especially low-rank adaptation (LoRA; Hu et al., 2021), have been shown to enable fine-tuning with only a fraction of the trainable parameters. But even when fine-tuning with LoRA matches the performance of full fine-tuning, are the solutions learned by the two methods really equivalent?
While full fine-tuning treats every parameter as trainable, LoRA treats the learned update to a weight matrix as the product of two low-rank matrices (Hu et al., 2021; Dettmers et al., 2023). While this parameterization is empirically effective, a principled explanation of the mechanism by which it matches the full fine-tuning performance has remained elusive. One explanation is offered by the intrinsic dimension hypothesis (Li et al., 2018; Aghajanyan et al., 2021), which posits that the update learned via fine-tuning has an intrinsically low intrinsic rank, suggesting that LoRA might recover an approximately equivalent solution to full fine-tuning. However, prior work has observed differences in the ability of LoRA and full fine-tuning to independently change the angle and magnitude with which a neuron transforms its input (Liu et al., 2024). Moreover, other work has also observed that LoRA has difficulty matching the performance of full fine-tuning on harder tasks, like code generation (Biderman et al., 2024; Zhuo et al., 2024) and long-form text generation (Ivison et al., 2023). Therefore, it is unclear if these findings indicate a limit in LoRA’s ability to fit to a specific downstream task, or if these methods learn inherently different solutions.
@@
Figure 2: Characterizing structural differences between solutions learnt by LoRA Vs full Fine-tuning. a) We measure the changes to the SVD of the pre-trained weights made during fine-tuning. We observe intruder dimensions introduced by LoRA in top ranking singular vectors but by full fine-tuning. b) Comparing a matrix fine-tuned with full fine-tuning or LoRA. c) Comparing a normal singular vs an intruder dimension to all pre-trained singular vectors.
@@
In this paper, we show that full fine-tuning and LoRA learn different solutions with characteristic differences in their spectral properties (as shown in Fig. 1) and different generalization behaviors outside the target task distribution. We observe:
@@
1. LoRA and full fine-tuning produce structurally different parameter updates, characterized by the existence of intruder dimensions. These are singular vectors, with large associated singular values, that are approximately orthogonal to the singular vectors in a pre-trained weight matrix. In contrast, fully fine-tuned models remain spectrally similar to the pre-trained model and do not contain intruder dimensions.
2. Behaviorally, LoRA fine-tuned models with intruder dimensions forget more of the pre-training distribution and exhibit less robust continual learning compared to full fine-tuning: LoRA fine-tuned models with intruder dimensions are inferior to fully fine-tuned models outside the adaptation task’s distribution, despite matching accuracy in distribution. However, higher-rank LoRA fine-tuned models, with identical adaptation task performance, more closely resemble fully fine-tuned models on these measures. Very high rank LoRA models, for e.g., full-rank LoRA, too forget more of their pre-training distribution—highlighting the fact that LoRA is not exempt from the general tradeoff between expressive power and generalization.
3. Even when a low-rank LoRA performs well on a target task, a higher-rank parameterization may still be preferable. While we observe that our low-rank LoRAs ($r\leq 8$) fit our downstream task distribution as well as full fine-tuning and high-rank LoRAs, using a high-rank ($r=64$) leads to models that both exhibit better generalization and robust adaptability. However, in order to take advantage of higher ranks, the LoRA updated models must be rank-stabilized (Kalajdzievski, 2023).
@@ 2 Background & Related Work
Methods for fine-tuning.
Pre-trained language models offer a foundation for downstream applications, eliminating the need to train from scratch (Ouyang et al., 2022; Devlin et al., 2019). Full fine-tuning, in which every parameter of a pre-trained model is updated, has been used for adaptation (Devlin et al., 2019; Liu et al., 2019). Low Rank Adaptation (LoRA; Hu et al., 2021), which represents the update to the weights as a product of two low-rank matrices, reduces computation and memory requirements relative to full fine-tuning. Past work has shown that LoRA matches full fine-tuning performance for tasks like sequence classification (Hu et al., 2021), instruction tuning (Dettmers et al., 2023; Ghosh et al., 2024), and chat (Dettmers et al., 2023). Other work has shown a gap in the performance of full fine-tuning and LoRA on harder tasks like code generation (Biderman et al., 2024; Zhuo et al., 2024). While we focus on models trained to similar accuracy, our observations of structural differences apply even to cases where LoRA does not fit to the adaptation task as well as full fine-tuning.
LoRA, formally.
Given a pre-trained weight matrix $W_{0}\in\mathbb{R}^{m\times n}$, full fine-tuning treats the learned matrix update as $\Delta W\in\mathbb{R}^{m\times n}$. Instead, LoRA decomposes $\Delta W$ into a product of two matrices such that $\Delta W=BA$, where $B\in\mathbb{R}^{m\times r}$, $A\in\mathbb{R}^{r\times n}$, and where the rank $r$ is generally $r\ll min(m,n)$. During prediction,
@@
$B$ is initialized to zero, and $A$ sampled from an isotropic Gaussian. All parameters in $B$ and $A$ are trained. From this we can see that while full fine-tuning mas $mn$ trainable parameters per weight matrix, LoRA only has $mr+rn$ trainable parameters. See Appendix D for derivation of gradients.
Figure 3: Cosine similarities between sorted singular vectors in the fine-tuned models to pre-trained models. (Right) Matrices fine-tuned with LoRA have a shift in singular vectors, as shown by blank columns, due to intruder dimensions (which are dissimilar to the pre-trained singular vectors). (Left) However, no such shift is found in the case of models trained via full fine-tuning.
LoRA Variants.
@@
Many variations of LoRA exist. Methods improve LoRA’s performance or memory-efficiency by initializing with the principal components of the underlying weight matrix (Meng et al., 2024), training with quantization (Dettmers et al., 2023), adaptively allocating different ranks (Zhang et al., 2023), or sequentially training multiple LoRAs (Xia et al., 2024). Other methods propose similar but alternative architectures (Liu et al., 2024; Kopiczko et al., 2024; Koohpayegani et al., 2024). Here, we focus on the original LoRA setup, as described in Hu et al. (2021). We leave a rigorous analysis of these variations and their impacts on our findings to future work.
Empirically, setting $\alpha=2r$ has been shown to improve results for higher ranks (Biderman et al., 2024) and is theoretically well motivated. (Kalajdzievski, 2023). We adopt this parameterization for most experiments in our paper.
Analysis of Solutions.
@@
Introduced by Li et al. (2018), the intrinsic dimension measure was used by Aghajanyan et al. (2021) to argue that the fine-tuning update for a pre-trained LLM has low intrinsic rank, explaining why only a small number of trainable parameters are necessary to reach 90% of full fine-tuning performance. This finding motivated Hu et al. (2021) to hypothesize that LoRA works because solutions of low intrinsic rank exist. But to our knowledge, no past work has compared the rank (or other properties of weight matrices) between LoRA and full-fine tuning on tasks where they are matched in performance. While Liu et al. (2024) showed that LoRA has difficulty changing directional and magnitude components of a neuron independently, while full fine-tuning does not, it is unclear if this difference is due to an inability of LoRA to fit as well as full fine-tuning to the adaptation task.
Recent work comparing LoRA to full fine-tuning has found that LoRA forgets less on previously learned information (Biderman et al., 2024) and more closely resembles the pre-trained model (Ghosh et al., 2024). Surprisingly, some experiments in the current study show opposite trends. However, there are significant differences in the datasets used for evaluation—(Biderman et al., 2024) investigated instruction tuning for language generation, while we mainly study sequence labeling tasks. Importantly, Biderman et al. (2024) study conditions when LoRA fine-tuned models fail to fit the adaptation task as well as full-finetuned models, and as a result also forget less of the pre-training distribution. However, we study models where the LoRA achieves the same performance as full fine-tuning, comparing generalization behavior at a fixed target task accuracy.
Singular Value Decomposition.
@@
The SVD decomposes a matrix $M\in\mathbb{R}^{m\times n}$ such that $M=U\Sigma V^{T}$, where $U\in\mathbb{R}^{m\times m}$ and $V\in\mathbb{R}^{n\times n}$ have orthonormal columns representing the singular vectors of $M$ and $\Sigma\in\mathbb{R}^{m\times n}$ is a diagonal matrix containing the singular values of $M$. $U$ and $V^{T}$ represent rotations that matrix $M$ performs, while $\Sigma$ represents scaling along those axes. Importantly, singular vectors ranked in order by their associated singular value capture the order of most important axes of transformation that the matrix performs.
3 Model Differences Between LoRA and Full Fine-Tuning
@@
Inspired by Sharma et al. (2024)’s findings that the Singular Value Decomposition (SVD, Klema & Laub, 1980) can be used to selectively prune singular vectors to improve model performance, this paper adopts the SVD of neural network parameters as a lens for understanding the changes that fine-tuning makes to pre-trained weights. Understanding how these dimensions change can give us insight into how a particular fine-tuning method changes the pre-trained model. In particular, we measure how well singular vectors in weight matrices fine-tuned with LoRA or full fine-tuning map to singular vectors in the pre-trained weights using their cosine similarity. These relationships are shown in Fig. 1 and Fig. 3, with color representing cosine similarity between pre-trained and fine-tuned singular vectors.
@@
Visually, we observe in Fig. 2(b) that LoRA and full fine-tuning’s singular vectors have very different similarities to the pre-trained singular vectors: singular vectors of models fine-tuned with LoRA appear to have, on average, much lower cosine similarity to pre-trained singular vectors in comparison to full fine-tuning. Interestingly, in LoRA fine-tuned models, we also observe the presence of high ranking singular vectors with very low cosine similarity to any pre-trained singular vector. [^1] Recall that in high dimensions, a vector can have low cosine similarity to a set of orthogonal vectors that span a space; see Appendix C for discussion. In Fig. 2(c), we show the difference between these vectors with low cosine similarity to the pre-trained singular vectors and normal singular vectors from the fine-tuned weights. This “new” dimension can be seen in Fig. 2(b) as the lone red dot in the bottom left corner. We name these “new” dimensions intruder dimensions, which we define formally as follows:
Definition 1
@@
A singular vector $y_{j}$ from the fine-tuned weight matrix $W_{tuned}$ is an intruder dimension if and only if $\max_{i}(cos(y_{j},x_{i}))<\epsilon$, where $\epsilon$ is a similarity threshold and $x_{i}$ is a singular vector in $W_{0}$.
@@
Examples of intruder dimensions may be seen in Fig. 3. Here, we plot the similarity matrix between the top 10 singular vectors (ranked by singular value) in the pre-trained and fine-tuned matrices. While full fine-tuning appears to have a clear one-to-one mapping, LoRA appears to have its mapping shifted by “blank” columns: these are intruder dimensions, with low cosine similarity to every pre-trained singular vector.
@@
Setup. We study RoBERTa-base (Liu et al., 2019), a pre-trained encoder-only language model, fine-tuned on six different sequence classification tasks. We train these models to similar performance on their respective downstream tasks to study how, at a similar level of performance, fully fine-tuned and LoRA fine-tuned models differ. See Appendix A for more fine-tuning details. We compute the total number of intruder dimensions across these models.
Algorithm: Finding Intruder Dimensions. Pre-trained weights $W_{0}$, fine-tuned weights $W_{tuned}$, cosine similarity threshold $\epsilon$, and number of fine-tuned singular vectors to examine $k$. $[U_{0},\Sigma_{0},V_{0}^{T}]\leftarrow\text{SVD}(W_{0})$ $[U_{\text{tuned}},\Sigma_{\text{tuned}},V_{\text{tuned}}^{T}]\leftarrow\text{SVD}(W_{\text{tuned}})$ num_intruders $\leftarrow 0$ for $j\leftarrow 1$ to $k$ do      $n\leftarrow$ of pre-trained singular vectors      if $\forall i\in\{1,\ldots,n\},\cos(U_{0}[i],U_{tuned}[j])<\epsilon$ then          num_intruders $\leftarrow$ num_intruders $+1$      end if end for return num_intruders
Figure 4: Outline of the procedure used to compute the total number of intruder dimensions introduced in a model.
LoRA fine-tuned models contain high-ranking intruder dimensions while fully fine-tuned models do not. To quantify the size of the set of intruder dimensions for a specific weight matrix, we use the algorithm described in Fig. 4. Concretely, we first compute the SVD of both the pre-trained and resulting LoRA and full fine-tuned weights. Following that, for each of the top $k$ highest-ranking singular vectors, we measure its maximum cosine similarity with all of the pre-trained singular vectors. If this maximum cosine similarity is less than some threshold $\epsilon$, we classify this singular vector as an intruder dimension. Note that both $k$, the number of fine-tuned singular vectors to examine, and $\epsilon$, the cosine similarity threshold, are hyperparameters; we verify the robustness of our findings for a wide range of $\epsilon$ and $k$ values in Fig. 5 and Fig. 11 respectively. To determine the number of intruder dimensions in a specific model, we run this algorithm for each weight matrix in the model and sum the total.
To characterize the differences in fine-tuning methods, we first evaluate the differences in the total number of intruder dimensions in the top 10 highest-ranking singular vectors ($k=10$). We repeat this procedure for a the range of $\epsilon$ values, our cosine similarity threshold. The results are presented in Fig. 5(a). We find that models trained with LoRA consistently contain intruder dimensions when their rank $r\leq 16$, particularly for low values of $\epsilon$. Interestingly, we observe that fully fine-tuned models almost never contain intruder dimensions in their top 10 singular vectors for epsilon values of about 0.6 to 0.9 across different settings. This means that full fine-tuning makes smaller changes to the same set of high contribution pre-trained singular vectors. Importantly, the number of intruder dimensions appears to drop as rank increases past a certain threshold, suggesting that the low-rank nature, as well as the update rule of LoRA, induces them to occur.
@@
Intruder dimensions exist even in tasks where LoRA fine-tuned models learn less than full fine-tuning. To test the validity of our findings for larger models and harder tasks, we study LLaMA-7B (Touvron et al., 2023a) and LLaMA2-7B (Touvron et al., 2023b) models fine-tuned on various instruction following datasets. These span math, code, and chat, and are considerably harder than our sequence classification tasks. See Appendix 3 for more details about these models.
Looking at Figs. 5(b), 5(c), and 5(d), we can clearly see intruder dimensions in the set of high ranking singular vectors for LoRA, even with a rank as high as $r=256$. Importantly, the $r=2048$ case of MetaMath does not have intruder dimensions and instead has a very similar curve to full fine-tuning. This supports the earlier finding that, as rank increases past a threshold, intruder dimensions disappear and LoRA begin to resemble full fine-tuning.
Interestingly, the full fine-tuned Magicoder model also has intruder dimensions for higher values of $\epsilon$. This is likely because, as mentioned by Biderman et al. (2024), there is a larger domain shift between coding tasks and the pre-training data in comparison to other natural language tasks. This difference likely causes full fine-tuning to make more aggressive changes to the model. But even in this case, LoRA models have many more intruder dimensions in their top 10 singular vectors than full fine-tuning (see Fig. 1).
@@
(a) Number of intruder dimensions in RoBERTa models fine-tuned on 6 different tasks.
(b) LLaMA-7B fine-tuned on Alpaca.
(c) LLaMA2-7B fine-tuned on MetaMathQA.
(d) LLaMA2-7B fine-tuned on Magicoder-Evol-Instruct.

### v2

@@ Abstract
Fine-tuning is a crucial paradigm for adapting pre-trained large language models to downstream tasks. Recently, methods like Low-Rank Adaptation (LoRA) have been shown to effectively fine-tune LLMs with an extreme reduction in trainable parameters. But, are their learned solutions really equivalent? We study how LoRA and full-finetuning change pre-trained models by analyzing the model’s weight matrices through the lens of their spectral properties. We find that LoRA and full fine-tuning yield weight matrices whose singular value decompositions exhibit very different structure: weight matrices trained with LoRA have new, high-ranking singular vectors, which we call intruder dimensions, while those trained with full fine-tuning do not. Further, we extend the finding that LoRA forgets less than full fine-tuning and find its forgetting is vastly localized to the intruder dimension – by causally intervening on the intruder dimensions by changing their associated singular values post-fine-tuning, we show that they cause forgetting. Moreover, scaling them down significantly improves modeling of the pre-training distribution with a minimal drop in downstream task performance. Given this, we should expect accumulating intruder dimensions to be harmful and lead to more forgetting. This will be amplified during continual learning because of sequentially fine-tuning, and we show that LoRA models do accumulate intruder dimensions here tend to perform worse in this setting, emphasizing the practicality of our findings.
@@ 1 Introduction
Figure 1: LoRA and full fine-tuning update the parameter space differently. Similarity matricies of pre- and post-fine-tuning singular vectors for LLaMA2-7B that characterize the spectral differences introduced during fine-tuning. Full fine-tuning retains most of the pre-training structure, while LoRA has a diagonal shift. Color shows cosine similarity.
@@
Adapting large, pre-trained models to downstream tasks via fine-tuning is a computation- and data-efficient way to create domain-specific models for a variety of tasks. The simplest approach is to fine-tune all parameters of the pre-trained model on downstream task data (Devlin et al., 2019; Ouyang et al., 2022). However, as pre-trained models grow larger, full fine-tuning becomes increasingly challenging and expensive. Recently, parameter-efficient fine-tuning (PEFT) methods, especially low-rank adaptation (LoRA; Hu et al., 2021), have been shown to enable fine-tuning with only a fraction of the trainable parameters. While LoRA can match full fine-tuning performance, are the solutions learned by the two methods similar?
@@
Figure 2: Characterizing structural differences between solutions learnt by LoRA & Full Fine-tuning. a) We measure the changes to the SVD of the pre-trained weights made during fine-tuning. We observe intruder dimensions introduced by LoRA in top ranking singular vectors but not by full fine-tuning. b) Comparing a matrix fine-tuned with full fine-tuning or LoRA. c) The intruder dimension shows near-zero absolute cosine similarity with all pre-trained singular vectors, in contrast to other singular vectors of the finetuned matrix.
While full fine-tuning treats every parameter as trainable, LoRA treats the learned update to a weight matrix as the product of two low-rank matrices (Hu et al., 2021). While this parameterization is empirically effective, a principled explanation of the mechanism by which it matches the full fine-tuning performance has remained elusive. One explanation is offered by the intrinsic dimension hypothesis (Li et al., 2018; Aghajanyan et al., 2021), which posits that the update learned via fine-tuning has a low intrinsic rank, suggesting that LoRA might recover an approximately equivalent solution to full fine-tuning. However, prior work has observed differences in the ability of LoRA and full fine-tuning to independently change the angle and magnitude with which a neuron transforms its input (Liu et al., 2024). Moreover, other work has also observed that LoRA has difficulty matching the performance of full fine-tuning on difficult tasks like code generation (Biderman et al., 2024; Zhuo et al., 2024) and long-form text generation (Ivison et al., 2023). Therefore, it is unclear if these findings indicate a limit in LoRA’s ability to fit to a specific downstream task, or if these methods learn inherently different solutions.
@@
In this paper, we show that full fine-tuning and LoRA learn different solutions with characteristic differences in their spectral properties (as shown in Fig. 1 for LLaMA2-7B (Touvron et al., 2023b)) and that these spectral differences are causally related to different model behaviors. We observe:
@@
1. LoRA and full fine-tuning produce structurally different parameter updates, characterized by the existence of intruder dimensions in weight matrices tuned by LoRA. Intruder dimensions are singular vectors with large associated singular values that are very dissimilar to the singular vectors in the pre-trained weight matrix. In contrast, fully fine-tuned models remain spectrally similar to the pre-trained model and do not contain intruder dimensions.
2. LoRA forgets less than full fine-tuning…but not always. We extend the findings of Biderman et al. (2024) that LoRA forgets less to the case even when there is equal fine-tuning performance between LoRA and full fine-tuning, showing that a difference in fit is not simply the cause of this finding but rather is inherent to these methods. However, this is not always the case: despite nearly identical fine-tuning task accuracies, we show that different selections of LoRA alpha and learning rate lead to starkly different generalization behaviors, even leading to LoRA forgetting more than full fine-tuning. We also find that models with the best generalization for each of these hyperparameter settings have the fewest intruder dimensions.
3. Intruder dimensions cause forgetting of the pre-training distribution. Scaling down the associated singular values of high-ranking intruder dimensions leads to a large drop in loss on the pre-training distribution (forgetting) but only a minimal drop in test performance. The drop in forgetting we observe when scaling down singular vectors is unique to intruder dimensions and indicates that they interfere with the pre-trained language modeling ability of these models. Given this finding, we should expect accumulating intruder dimensions to be harmful and lead to more forgetting. To amplify this accumulation and examine its effect, we fine-tune in a continual learning setting (sequentially fine-tuning on many tasks) and show that LoRA models do indeed tend to forget more on previously learned tasks in this setting, providing additional support for our findings.
@@ 2 Background & Related Work
Methods for fine-tuning. Pre-trained language models offer a foundation for downstream applications, eliminating the need to train from scratch (Ouyang et al., 2022; Devlin et al., 2019). Full fine-tuning, in which every parameter of a pre-trained model is updated, is commonly used (Devlin et al., 2019; Liu et al., 2019). Low Rank Adaptation (LoRA; Hu et al., 2021), which represents the update to the weights as a product of two low-rank matrices, reduces computation and memory requirements relative to full fine-tuning. Past work has shown that LoRA matches full fine-tuning performance for tasks like sequence classification (Hu et al., 2021), instruction tuning (Dettmers et al., 2023; Ghosh et al., 2024), and chat (Dettmers et al., 2023). Other work has shown a gap in performance on harder tasks like code generation (Biderman et al., 2024; Zhuo et al., 2024). We focus our investigation on both cases to ensure our findings generalize to all use cases.
LoRA, formally. Given a pre-trained weight matrix $W_{0}\in\mathbb{R}^{m\times n}$, full fine-tuning treats the learned matrix update as $\Delta W\in\mathbb{R}^{m\times n}$. Instead, LoRA decomposes $\Delta W$ into a product of two matrices such that $\Delta W=BA$, where $B\in\mathbb{R}^{m\times r}$, $A\in\mathbb{R}^{r\times n}$, and where the rank $r$ is generally $r\ll min(m,n)$. During prediction,
@@
$B$ is initialized to zero, and $A$ sampled from an isotropic Gaussian. All parameters in $B$ and $A$ are trained. From this we can see that while full fine-tuning has $mn$ trainable parameters per weight matrix, LoRA only has $mr+rn$. See Appendix E for derivation of LoRA adapter gradients.
@@
LoRA Variants. Many variations of LoRA exist. Methods improve LoRA’s performance or memory-efficiency by initializing with the principal (Meng et al., 2024) or minor (Wang et al., 2024) components of the underlying weight matrix, training with quantization (Dettmers et al., 2023), adaptively allocating different ranks (Zhang et al., 2023), or sequentially training multiple LoRAs (Xia et al., 2024). Other methods propose similar but alternative architectures (Liu et al., 2024; Kopiczko et al., 2024; Koohpayegani et al., 2024). Other work has also proposed low rank manipulations to the activations instead of the weights (Wu et al., 2024). Although the primary focus of our study is on the original LoRA setup (Hu et al., 2021), we also study a few LoRA variants (Appendix O). While we leave a rigorous analysis of all possible variants to future work, our preliminary experiments show that our findings generalize to several variants. Additionally, we also demonstrate the robustness of our findings across a range of LoRA hyperparameter settings (Appendices A.4, M, N).
@@
Analysis of Solutions. The intrinsic dimension measure (Li et al., 2018) was used by Aghajanyan et al. (2021) to argue that the fine-tuning update for a pre-trained LLM has low intrinsic rank, explaining why only a small number of trainable parameters are necessary to reach 90% of full fine-tuning performance. This finding motivated Hu et al. (2021) to hypothesize that LoRA works because solutions of low intrinsic rank exist. But to our knowledge, no past work has compared the rank (or other properties of weight matrices) between LoRA and full-fine tuning on tasks where they are matched in performance. While Liu et al. (2024) showed that LoRA has difficulty changing directional and magnitude components of a neuron independently, it is unclear if this difference is due to an inability of LoRA to fit as well as full fine-tuning to the adaptation task.
Relation to Biderman et al. (2024). Recent work comparing LoRA to full fine-tuning has found that LoRA forgets less when fine-tuned on math and code (Biderman et al., 2024) and more closely resembles the pre-trained model (Ghosh et al., 2024). We extend the findings of Biderman et al. (2024) to the case when there is equal fine-tuning performance between LoRA and full fine-tuning, showing that a difference in fit to the fine-tuning task is not simply the cause of this finding but rather is inherent to these methods.
@@
Singular Value Decomposition. The SVD decomposes a matrix $M\in\mathbb{R}^{m\times n}$ such that $M=U\Sigma V^{T}$, where $U\in\mathbb{R}^{m\times m}$ and $V\in\mathbb{R}^{n\times n}$ have orthonormal columns representing the singular vectors of $M$ and $\Sigma\in\mathbb{R}^{m\times n}$ is a diagonal matrix containing the singular values of $M$. $U$ and $V^{T}$ represent rotations that matrix $M$ performs, while $\Sigma$ represents scaling along those axes. Singular vectors, ordered by singular values, reveal a matrix’s most important axes of transformation.
3 Structural Differences
Figure 3: LoRA and full fine-tuning learn distinct structural solutions. LoRA introduces intruder dimensions (represented by outlined columns).
@@ 3 Structural Differences
Inspired by Sharma et al. (2024)’s finding that the singular value decomposition (SVD, Klema and Laub, 1980) can be used to selectively prune singular vectors to improve model performance, this paper adopts the SVD of neural network parameters as a lens for understanding the changes that fine-tuning makes to pre-trained weights. Understanding how these dimensions change can give us insight into how a particular fine-tuning method changes the pre-trained model. In particular, we study how well singular vectors in weight matrices fine-tuned with LoRA or full fine-tuning map to singular vectors in the pre-trained weights (using cosine similarity).
@@
Visually, we observe in Fig. 2(b) that LoRA and full fine-tuning’s singular vectors have very different similarities to the pre-trained singular vectors: singular vectors of models fine-tuned with LoRA appear to have, on average, much lower cosine similarity to pre-trained singular vectors in comparison to full fine-tuning. Interestingly, in LoRA fine-tuned models, we also observe the presence of high ranking singular vectors with very low cosine similarity to any pre-trained singular vector. [^1] Recall that in high dimensions, a vector can have low cosine similarity to a set of orthogonal vectors that span a space; see Appendix D for discussion. In Fig. 2(c), we show the difference between these vectors with low cosine similarity to the pre-trained singular vectors and normal singular vectors from the fine-tuned weights. This “new” dimension can be seen in Fig. 2(b) as the lone red dot in the bottom left corner. We name these “new” dimensions intruder dimensions, which we define formally as follows:
Definition 3.1.
@@
A singular vector $y_{j}$ from the fine-tuned weight matrix $W_{tuned}$ is an intruder dimension if and only if $\max_{i}(cos(y_{j},x_{i}))<\epsilon$, where $\epsilon$ is a similarity threshold and $x_{i}$ are the singular vectors of $W_{0}$.
Algorithm 1 Finding intruder dimensions.
1:Pre-trained weights $W_{0}$, fine-tuned weights $W_{\text{t}}$, cosine similarity threshold $\epsilon$, # of fine-tuned singular vectors to examine $k$.
2:$[U_{0},\Sigma_{0},V_{0}^{\!\top}]\leftarrow\text{SVD}(W_{0})$
3:$[U_{t},\Sigma_{t},V_{t}^{\!\top}]\leftarrow\text{SVD}(W_{\text{tuned}})$
4:${{\rm n\_intruders}}\leftarrow 0$
5:$n\leftarrow$ of pre-trained singular vectors
6:for $j\leftarrow 1$ to $k$ do
7:     if $\forall i\leq n:\,\cos\!\bigl{(}U_{0}[i],\,U_{t}[j]\bigr{)}<\epsilon$ then
8:         ${{\rm n\_intruders}}\leftarrow{{\rm n\_intruders}}+1$
9:     end if
10:end for
11:return ${{\rm n\_intruders}}$
@@
Examples of intruder dimensions may be seen in Fig. 3. Here, we plot the similarity matrix between the top 10 singular vectors (ranked by singular value) in the pre-trained and fine-tuned matrices. While full fine-tuning appears to have a clear one-to-one mapping, LoRA appears to have its mapping shifted by “blank” columns (outlined in magenta): these are intruder dimensions, with low cosine similarity to every pre-trained singular vector. A zoomed out version of this plot can be seen in Fig. 1, in which we see an off diagonal shift due to intruder dimensions.
@@
Our Models. We study LLaMA2-7B (Touvron et al., 2023b) and RoBERTa-base (Liu et al., 2019). RoBERTa-base is a pre-trained encoder-only language model and we fine-tune it on six different sequence classification tasks. See Appendix B.3 for fine-tuning details. LLaMA2-7B is a pre-trained decoder-only language model, and we study it when fine-tuned on either code or math. These checkpoints are provided by Biderman et al. (2024). We also study LLaMA-7B (Touvron et al., 2023a) models fine-tuned on instruction following. See Appendix K for more details about these models. Importantly, these LLaMA models span math, code, and chat, which are considerably harder than sequence classification tasks. This ensures wide coverage of LoRA use cases.
(a) LLaMA-7B fine-tuned on Alpaca.
(b) LLaMA2-7B fine-tuned on MetaMathQA.
(c) LLaMA2-7B fine-tuned on Magicoder-Evol-Instruct.
@@
(d) Number of intruder dimensions in RoBERTa models fine-tuned on 6 different tasks.